A satellite is put into an elliptical orbit around
Answer #1

According to law of conservation of angular momentum, \frac{v_{1}^{2}}{r_{1}} = \frac{v_{2}^{2}}{r_{2}} aroud a planet

Here r1= 6400+225 =6625km

and v1 = 9.85km/s

therefore r2 = 6400+ha , here 6400km is the radius of the earth

h_{a} = v_{2}^{2}\frac{r_{1}}{v_{1}^{2}}-6400km

h_{a} = v_{2}^{2}X6625/9.85^{2}-6400

NOTE: in the given data the velocity at apogee is missing, when you substitue we get ha.